25 ml AgNO3 reacts with 5 mM NaCl; 35 ml of the silver solution is required to titrate a sample of ZnCl2∙6H2O. How many milligrams of the hydrate are present in the sample?
1
Expert's answer
2020-05-15T10:05:35-0400

ZnCl2 + 2AgNO3 = Zn(NO3)2 + 2AgCl

n(AgNO3) = V(AgNO3) ∙ C(AgNO3) = 0.035 / (V(NaCl) ∙ 0.005 / 0.025) = 1.75 / V (NaCl)

m(ZnCl2 ∙ 6H2O) = 2 ∙ n(AgNO3) ∙ M(ZnCl2 ∙ 6H2O) = 488.6 ∙ n(AgNO3) = 279.2 ∙ V(NaCl)